# FC-KAN: Function Combinations in Kolmogorov-Arnold Networks

### v1

@@ Abstract
In this paper, we introduce FC-KAN, a Kolmogorov-Arnold Network (KAN) that leverages combinations of popular mathematical functions such as B-splines, wavelets, and radial basis functions on low-dimensional data through element-wise operations. We explore several methods for combining the outputs of these functions, including sum, element-wise product, the addition of sum and element-wise product, quadratic function representation, and concatenation. In our experiments, we compare FC-KAN with multi-layer perceptron network (MLP) and other existing KANs, such as BSRBF-KAN, EfficientKAN, FastKAN, and FasterKAN, on the MNIST and Fashion-MNIST datasets. A variant of FC-KAN, which uses a combination of outputs from B-splines and Difference of Gaussians (DoG) in the form of a quadratic function, outperformed all other models on the average of 5 independent training runs. We expect that FC-KAN can leverage function combinations to design future KANs. Our repository is publicly available at: https://github.com/hoangthangta/FC_KAN.
@@ 1 Introduction
Recent work by Liu et al. [1, 2] highlights the use of learnable activation functions as "edges" to fit training data better, in contrast to the traditional use of fixed activation functions as "nodes" in multi-layer perceptrons (MLPs). The foundation of KANs is based on the Kolmogorov-Arnold representation theorem (KART), which asserts that any continuous function of multiple variables can be represented as a sum of continuous functions of a single variable [3]. Motivated by this idea, numerous researchers have worked on creating different types of KANs utilizing widely known polynomials and basis functions. While most of these works use a single function in constructing KANs [4, 5, 6, 7], there are a few works using function combinations in KANs’ structure. Ta [8] designed BSRBF-KAN, by combining B-splines and Gaussian Radial Basis Functions within each network layer using data additions. However, this network encounters a memory issue when performing element-wise operations in high-dimensional data, restricting the use of other element-wise operations, such as element-wise product (Hadamard product). In another work, Yang et al. [9] leveraged function combinations to develop optimal activation functions at each node through an adaptive strategy, overcoming the limitations of single activation functions in their S-KAN and S-ConvKAN models.
@@
In this paper, we propose a novel KAN, FC-KAN (Function Combinations in Kolmogorov-Arnold Networks), which leverages various functions to capture data features throughout the network layers and combines them in low-dimensional spaces, such as the output layer, using various methods mainly based on element-wise operations, including sum, product, the combination of sum and product, quadratic function representation, and concatenation. We avoid using a 3rd-degree function due to its higher computational demands on data tensors, which can lead to memory errors. As a result, the model is able to capture more data features, leading to improved performance on the MNIST and Fashion-MNIST datasets compared to other KAN networks. Additionally, employing n-degree functions aligns with the core concept of KAN, where they are used both to capture input data and to represent data features in the output. We expect this exciting development may lead to the proliferation of basis function combinations in neural networks.
@@
Aside from this section, the paper is organized as follows: Section 2 discusses related work on KART and KANs. Section 3 details our methodology, covering KART, the design of the KAN architecture, several existing KANs, and FC-KAN. Section 4 presents our experiments, comparing FC-KAN variants with MLP and other KANs using data from the MNIST and Fashion-MNIST datasets. Additionally, this section includes a comparison of combination types within FC-KAN. We address some limitations of our study in Section 5. Lastly, Section 6 offers our conclusions and potential directions for future research.
@@ 2 Related Works
Despite the long history of KART’s application in neural networks, it had not garnered significant attention in the research community until the recent study by Liu et al. [1]. They suggested moving away from strict adherence to KART and instead generalizing it to develop KANs with additional neurons and layers. Our intuition aligns with this perspective as it helps to mitigate the issue of non-smooth functions when applying KART to neural networks. Consequently, KANs have the potential to outperform MLPs in both accuracy and interpretability for small-scale AI + Science tasks. However, the resurgence of KANs has also faced criticism from Dhiman [20], who argue that KANs are essentially MLPs with spline-based activation functions, in contrast to traditional MLPs with fixed activation functions.
@@
KANs have introduced a new perspective to the scientific community since May 2024. Many KAN designs utilize well-known mathematical functions, particularly those capable of handling curves, such as B-Splines [21] (Original KAN [1], EfficientKAN [^1] https://github.com/Blealtan/efficient-kan, BSRBF-KAN [8]), Gaussian Radial Basis Functions (GRBFs) (FastKAN [4], DeepOKAN [22], BSRBF-KAN [8]), Reflection SWitch Activation Function (RSWAF) in FasterKAN [5], Chebyshev polynomials (TorchKAN [23], Chebyshev KAN [7]), Legendre polynomials (TorchKAN [23]), Fourier transform (FourierKAN [^2] https://github.com/GistNoesis/FourierKAN/, FourierKAN-GCF [24]), wavelets [6, 25], and other polynomial functions [26]. While these works focus on single functions to design KANs, Ta [8] mentioned the combination of functions – B-Splines and radial basis functions – in designing KANs. Their BSRBF-KAN showed better convergence on the training data for MNIST and Fashion-MNIST. In another work, Yang et al. [9] utilized function combinations to create optimal activation functions at each node using an adaptive strategy, addressing the drawbacks of single activation functions in their S-KAN model. They also extended S-KAN to S-ConvKAN, which showed superior performance in image classification tasks, outperforming CNNs and KANs with comparable structures.
@@ 3.1 Kolmogorov-Arnold Representation Theorem
A KAN is based on KART, which asserts that any continuous multivariate function $f$ defined on a bounded domain can be represented as a finite combination of continuous single-variable functions and their additions [27, 28]. For a set of variables $\mathbf{x}={x_{1},x_{2},\ldots,x_{n}}$, where $n$ is the number of variables, the multivariate continuous function $f(\mathbf{x})$ is expressed as:
@@ 3.3 Implementation of the current KANs
LiuKAN [^3] We refer to this as LiuKAN, following the first author’s last name [1], while another work [6] refers to it as Spl-KAN. was implemented by Liu et al. [1] by using the residual activation function $\phi(x)$ as the sum of the base function and the spline function with their corresponding weight matrices $w_{b}$ and $w_{s}$:
@@
where $b(x)$ equals to $silu(x)$ and $spline(x)$ is expressed as a linear combination of B-splines. Each activation function is activated with $w_{s}=1$ and $spline(x)\approx 0$, while $w_{b}$ is initialized by using Xavier initialization.
@@
FastKAN can speed up training over EfficientKAN by using GRBFs to approximate the 3-order B-spline and employing layer normalization to keep inputs within the RBFs’ domain [4]. These modifications simplify the implementation without sacrificing accuracy. The GRBF has the formula:
@@
where $r=\|x-c\|$ is the distance between an input vector $x$ and a center $c$, and $\epsilon$ ($epsilon>0$) is a sharp parameter that controls the width of the Gaussian function. FastKAN uses a special form of RBFs where $\epsilon=\frac{1}{2}$ as [4]:
@@
FasterKAN outperforms FastKAN in both forward and backward processing speeds [5]. It uses Reflectional Switch Activation Functions (RSWAFs), which are variants of RBFs and straightforward to compute due to their uniform grid structure. The RSWAF function is shown as:
@@ 3.4 FC-KAN
Ta [8] introduced the idea of combining functions, such as B-splines and GRBFs in BSRBF-KAN, to improve convergence when training models for image classification. However, their method was limited to element-wise addition of function outputs in each layer, without exploring other matrix operations like multiplications or different combinations. We argue that this approach might not effectively capture the input data’s features. It is important to note that multiplying matrices (or the Hadamard product) of high-dimensional data can lead to memory errors on GPU/CPU devices. Therefore, it is wise to perform these operations on low-dimensional data, such as the output layer of a neural network, in data classification problems.
@@
which $f_{i,l}$ is the function $f_{i}$ at the layer $l$. So we have a set of outputs $O=\{\mathbf{o_{1}},\mathbf{o_{2}},...,\mathbf{o_{n}}\}$ corresponding to the number of functions used. Then, we apply several methods to combine these outputs using element-wise operations, including the sum output (Equation 15a), the element-wise product output (Equation 15b), the sum and element-wise product output(Equation 15c), the quadratic function output (Equation 15d), and the concatenated output (Equation 15e). The final outputs from combination methods will maintain the same data dimension as each element $\mathbf{o}_{i}$ in $O$, except in the concatenation method, where the data size will be $\mathbf{o}_{i}$ multiplied by the number of outputs being combined.
$\mathbf{o}_{sum}=\sum_{i=1}^{n}\mathbf{o}_{i}$ (15a)
$\mathbf{o}_{prod}=\bigotimes_{i=1}^{n}\mathbf{o}_{i}=\mathbf{o}_{1}\odot\mathbf{o}_{2}\odot\cdots\odot\mathbf{o}_{n}$ (15b)
$\mathbf{o}_{sum+prod}=\sum_{i=1}^{n}\mathbf{o}_{i}+\bigotimes_{i=1}^{n}\mathbf{o}_{i}$ (15c)
$\begin{split}\mathbf{o}_{quad}=&\sum_{i=1}^{n}\mathbf{o}_{i}+\bigotimes_{i=1}^{n}\mathbf{o}_{i}+\mathbf{o}_{1}^{2}+\mathbf{o}_{2}^{2}+\dots+\mathbf{o}_{n}^{2}\\ =&\sum_{i=1}^{n}\mathbf{o}_{i}+\bigotimes_{i=1}^{n}\mathbf{o}_{i}+\mathbf{o}_{1}\odot\mathbf{o}_{1}+\mathbf{o}_{2}\odot\mathbf{o}_{2}+\dots+\mathbf{o}_{n}\odot\mathbf{o}_{n}\end{split}$ (15d)
$\mathbf{o}_{concat}=\bigoplus_{i=1}^{n}\mathbf{o}_{i}=\mathbf{o}_{1}\oplus\mathbf{o}_{2}\oplus\cdots\oplus\mathbf{o}_{n}$ (15e)
Output combinations can utilize functions of degree 3 or higher, but these may significantly increase computational complexity, especially in matrix multiplication. Additionally, using more functions results in a larger number of outputs, which can further complicate data combination calculations. To manage this complexity, we prefer to restrict output combinations to quadratic functions involving up to two outputs. For instance, to combine DoG and B-Splines at the output, we can use the following quadratic function formula:

### v2

@@ Abstract
In this paper, we introduce FC-KAN, a Kolmogorov-Arnold Network (KAN) that leverages combinations of popular mathematical functions such as B-splines, wavelets, and radial basis functions on low-dimensional data through element-wise operations. We explore several methods for combining the outputs of these functions, including sum, element-wise product, the addition of sum and element-wise product, representations of quadratic and cubic functions, concatenation, linear transformation of the concatenated output, and others. In our experiments, we compare FC-KAN with a multi-layer perceptron network (MLP) and other existing KANs, such as BSRBF-KAN, EfficientKAN, FastKAN, and FasterKAN, on the MNIST and Fashion-MNIST datasets. A variant of FC-KAN, which uses a combination of outputs from B-splines and Difference of Gaussians (DoG) in the form of a quadratic function, outperformed all other models on the average of 5 independent training runs. We expect that FC-KAN can leverage function combinations to design future KANs. Our repository is publicly available at: https://github.com/hoangthangta/FC_KAN.
@@ 1 Introduction
The work by Liu et al. [1, 2] helped to revive the usage of the Kolmogorov-Arnold representation theorem (KART) in neural networks. They highlighted the use of learnable activation functions as "edges" to fit training data better, in contrast to the traditional use of fixed activation functions as "nodes" in multi-layer perceptrons (MLPs). The foundation of KANs is based on KART, which was developed to address Hilbert’s 13th problem [3], specifically, the statement, “Prove that the equation of seventh degree $x^{7}+ax^{3}+bx^{2}+cx+1=0$ is not solvable by means of any continuous functions of only two variables”. KART asserts that any continuous function of multiple variables can be represented as a sum of continuous functions of a single variable [4]. Motivated by this idea, numerous researchers have worked on creating different types of KANs utilizing widely known polynomials and basis functions.
While these works use a single function (not mentioning activation functions) in constructing KANs [5, 6, 7, 8], there are a few ones using function combinations [9, 10, 11, 2]. In several works that designed MLP-based neural networks, the function combinations typically occur in activation functions to improve model performance and stability [12, 13, 14]. In KANs, the combinations occur directly on the functions used to fit input data, rather than through activation functions. For instance, BSRBF-KAN [9], which combines B-splines and Gaussian Radial Basis Functions (GRBFs) in all network layers. However, the authors only applied addition operations to data tensors without incorporating element-wise multiplications, which we believe does not effectively capture data features. Additionally, element-wise multiplications may lead to memory issues or slow calculations when performed on high-dimensional data. While Liu et al. [2] used element-wise multiplications alongside additions in MultKAN to enhance data capture, their focus was limited to small-scale examples.
@@
With the purpose of exploiting data features efficiently, we propose a novel KAN, FC-KAN (Function Combinations in Kolmogorov-Arnold Networks), which leverages various functions to capture input data throughout the network layers and combines them in low-dimensional spaces, such as the output layer, using various methods mainly based on element-wise operations, including sum, product, the combination of sum and product, representations of quadratic and cubic functions, and concatenation. We avoid using a higher-degree function due to their increased computational demands on data tensors, which can lead to memory errors. As a result, the model is able to capture more data features, leading to improved performance on the MNIST and Fashion-MNIST datasets compared to other KAN networks. Additionally, employing n-degree functions aligns with the core concept of KAN, where they are used both to capture input data and to represent data features in the output. We expect this exciting development to drive the proliferation of basis function combinations in neural networks. In summary, our main contributions are:
  • Construct FC-KAN to demonstrate the effectiveness of function combinations through various methods applied to low-dimensional data, such as model outputs.
  • Compare combination methods on the function outputs using mainly element-wise operations to search for proper methods.
@@
Aside from this section, the paper is organized as follows: Section 2 discusses related work on KART and KANs. Section 3 details our methodology, covering KART, the design of the KAN architecture, several existing KANs, and FC-KAN. Section 4 presents our experiments, comparing FC-KAN variants with MLP and other existing KANs using data from the MNIST and Fashion-MNIST datasets. Additionally, this section includes a comparison of combination methods within FC-KAN. We address some limitations of our study in Section 5. Lastly, Section 6 offers our conclusions and potential directions for future research.
@@ 2.1 KART and KAN
Despite the long history of KART’s application in neural networks, it had not garnered significant attention in the research community until the recent study by Liu et al. [1]. They suggested moving away from strict adherence to KART and generalizing it to develop KANs with additional neurons and layers. Our intuition aligns with this perspective as it helps to mitigate the issue of non-smooth functions when applying KART to neural networks. Consequently, KANs have the potential to outperform MLPs in both accuracy and interpretability for small-scale AI + Science tasks. However, the resurgence of KANs has also faced criticism from Dhiman [25], who argue that KANs are essentially MLPs with spline-based activation functions, in contrast to traditional MLPs with fixed activation functions.
@@
By introducing a new perspective to the scientific community in the new design of neural networks, KANs inspired many works to prove their effectiveness by topics, including expensive problems [26], keyword spotting [27], mechanics problems [28], quantum computing [29, 30, 31], survival analysis [32], time series forecasting [33, 34, 35, 36, 37], and vision tasks [38, 39, 40]. Also, many novel KANs utilize well-known mathematical functions, particularly those capable of handling curves, such as B-Splines [41] (Original KAN [1], EfficientKAN [^1] https://github.com/Blealtan/efficient-kan, BSRBF-KAN [9]), Gaussian Radial Basis Functions (GRBFs) (FastKAN [5], DeepOKAN [28], BSRBF-KAN [9]), Reflection SWitch Activation Function (RSWAF) in FasterKAN [6], Chebyshev polynomials (TorchKAN [42], Chebyshev KAN [8]), Legendre polynomials (TorchKAN [42]), Fourier transform (FourierKAN [^2] https://github.com/GistNoesis/FourierKAN/, FourierKAN-GCF [43]), wavelets [7, 44], and other polynomial functions [45].
2.2 Function combinations in KANs and other neural networks
Several works utilize the function combinations to design novel KANs. Ta [9] mentioned the combination of functions – B-Splines and radial basis functions – in designing KANs. Their BSRBF-KAN showed better convergence on the training data for MNIST and Fashion-MNIST. Liu et al. [2] introduced MultKAN, which consists of multiplication operations to capture multiplicative structures in data better. Unlike KAN [1], which directly copies nodes, MultKAN uses both addition nodes (copied from subnodes) and multiplication nodes (multiplying multiple subnodes). However, their work focused on small-scale examples only. Yang et al. [10] utilized function combinations to create optimal activation functions at each node using an adaptive strategy, addressing the drawbacks of single activation functions in their S-KAN model. They also extended S-KAN to S-ConvKAN, which showed superior performance in image classification tasks, outperforming CNNs and KANs with comparable structures. In another work,  Altarabichi [11] proposed the replacement of the sum with the average function in KAN neurons that can improve the model performance and keep the training stability in their DropKAN [46].
Before the appearance of KANs, several studies focused on constructing combinations of activation functions (along with their weights) or utilizing a set of flexible activations in neural networks to enhance model performance and stability. Jie et al. [12] introduced a new family of flexible activation functions for LSTM cells, along with another family developed by combining ReLUs and ELUs. Their findings demonstrate that LSTM models using the P-Sig-Ramp flexible activations significantly improve time series forecasting. Additionally, the P-E2-ReLU activation exhibits enhanced stability and performance in lossy image compression tasks with convolutional autoencoders. Xu and Chen [13] investigated a selection of widely used activation functions across various datasets in classification and regression tasks. They then created combinations of the best-performing activation functions within a convex restriction, showing improved performance compared to the corresponding base activation functions in a standard broad learning system. In another study, Zhang [14] developed a novel deep neural network (DNN) that optimizes activation function combinations for different neurons based on extensive simulations. The experiments showed that their DNNs outperformed those that rely on a single activation function. Instead of using the inner product between data vectors and weights,  Fan et al. [47] replaced this operation with a quadratic function in neurons within deep quadratic neural networks. Quadratic neurons provide enhanced expressive capability compared to conventional neurons, highlighting the advantages of quadratic networks in terms of expressive efficiency, unique representation, compact architecture, and computational capacity.
In addition to exploring function combinations in mathematics, it is also possible to merge function outputs using tensor operations. When the data dimensions are the same, element-wise operations such as summation or product can be applied to combine these outputs [48] which appear in many multi-modal tasks [49, 50, 51]. However, the computational challenges posed by high-dimensional data, such as inefficiencies in tensor-product operations, have led to a growing body of work aimed at optimizing and accelerating these processes [52, 53]. Furthermore, other studies have focused on tensor fusion and tensor decomposition techniques to extract meaningful features from data tensors and to simplify data combinations [54, 55, 56, 57].
@@ 3.1 Kolmogorov-Arnold Representation Theorem
A KAN is based on KART, which states that any continuous multivariate function $f$ defined on a bounded domain can be represented as a finite combination of continuous single-variable functions and their additions [58, 59]. For a set of variables $\mathbf{x}={x_{1},x_{2},\ldots,x_{n}}$, where $n$ is the number of variables, the multivariate continuous function $f(\mathbf{x})$ is expressed as:
@@ 3.3 Implementation of the current KANs
Figure 1: Left: The structure of KAN(2,3,1). Right: The simulation of how to calculate $\phi_{1,1,1}$ by control points and B-splines.
LiuKAN [^3] We refer to the original KAN as LiuKAN, following the first author’s last name [1], while another work [7] refers to it as Spl-KAN. was implemented by Liu et al. [1] by using the residual activation function $\phi(x)$ as the sum of the base function and the spline function with their corresponding weight matrices $w_{b}$ and $w_{s}$:
@@
$b(x)=silu(x)=\frac{x}{1+e^{-x}}$ (7)
$spline(x)=\sum_{i}c_{i}B_{i}(x)$ (8)
where $b(x)$ equals to $silu(x)$ as in Equation 7 and $spline(x)$ is expressed as a linear combination of B-splines $B_{i}$s and control points (coefficients) $c_{i}$s as in Equation 8. Each activation function is activated with $w_{s}=1$ and $spline(x)\approx 0$, while $w_{b}$ is initialized by using Xavier initialization. Figure 1 shows the structure of KAN(2,3,1) and how to calculate an inner function $\phi$ by control points and B-splines.
@@
FastKAN can speed up training over EfficientKAN by using GRBFs to approximate the 3-order B-spline and employing layer normalization to keep inputs within the RBFs’ domain [5]. These modifications simplify the implementation without sacrificing accuracy. The RBF has the formula:
@@
where $r=\|x-c\|$ is the distance between an input vector $x$ and a center $c$, and $\epsilon$ ($epsilon>0$) is a sharp parameter that controls the width of the Gaussian function. FastKAN uses a special form of RBFs, Gaussian RBFs where $\epsilon=\frac{1}{2}$ as [5]:
@@
FasterKAN outperforms FastKAN in both forward and backward processing speeds [6]. It uses Reflectional Switch Activation Functions (RSWAFs), which are variants of RBFs and straightforward to compute due to their uniform grid structure. The RSWAF function is shown as follows:
@@ 3.4 FC-KAN
Ta [9] introduced the idea of combining functions, such as B-splines and GRBFs in BSRBF-KAN, to improve convergence when training models for image classification. However, their method was limited to element-wise addition of function outputs in each layer, without exploring other matrix operations like multiplications or different combinations. We argue that this approach might not effectively capture the input data’s features. It is important to note that multiplying matrices of high-dimensional data can lead to memory errors or inefficient running time on GPU/CPU devices. Therefore, it is wise to perform these operations on low-dimensional data, such as the output layer of a neural network, in data classification problems.
Figure 2: The structure of FC-KAN and the three types of combined outputs: element-wise, concatenation, and linearization.
@@
which $f_{i,l}$ is the function $f_{i}$ at the layer $l$. So we have a set of outputs $O=\{\mathbf{o_{1}},\mathbf{o_{2}},...,\mathbf{o_{n}}\}$ corresponding to the number of functions used. Note that all outputs in $O$ have the same size.
After that, we combine function outputs to obtain the element-wise output, the concatenated output, and the linearized output, as shown in Figure 2. The element-wise output is formed by using element-wise operations, including the sum output (Equation 17a), the element-wise product output (Equation 17b), the sum and element-wise product output (Equation 17c), the quadratic and cubic function outputs (Equation 17d and Equation 17e). The concatenated output is formed by simply concatenating the function outputs together as in Equation 17f. The linearized output is formed by passing the concatenated output through a linear transformation (with matrix multiplication by $W$ and addition of bias $b$) as in Equation 17g. The element-wise output and the linearized output will maintain the same data dimension as each element $\mathbf{o_{i}}$ in $O$, except in the concatenated output, where the data size will be $\mathbf{o_{i}}$ multiplied by the number of function outputs being combined.
$\mathbf{o}_{sum}=\sum_{i=1}^{n}\mathbf{o}_{i}=\mathbf{o}_{1}+\mathbf{o}_{2}+\cdots+\mathbf{o}_{n}$ (17a)
$\mathbf{o}_{prod}=\bigodot_{i=1}^{n}\mathbf{o}_{i}=\mathbf{o}_{1}\odot\mathbf{o}_{2}\odot\cdots\odot\mathbf{o}_{n}$ (17b)
$\mathbf{o}_{sum+prod}=\mathbf{o}_{sum}+\mathbf{o}_{prod}=\sum_{i=1}^{n}\mathbf{o}_{i}+\bigodot_{i=1}^{n}\mathbf{o}_{i}$ (17c)
$\begin{split}\mathbf{o}_{quad}=&\mathbf{o}_{sum+prod}+\sum_{i=1}^{n}\mathbf{o_{i}}\odot\mathbf{o_{i}}\\ =&\mathbf{o}_{sum+prod}+\mathbf{o}_{1}\odot\mathbf{o}_{1}+\mathbf{o}_{2}\odot\mathbf{o}_{2}+\dots+\mathbf{o}_{n}\odot\mathbf{o}_{n}\end{split}$ (17d)
$\begin{split}\mathbf{o}_{cubic}=\mathbf{o}_{quad}\odot\mathbf{o}_{sum}\end{split}$ (17e)
$\mathbf{o}_{concat}=concat(\mathbf{o}_{1},\mathbf{o}_{2},\cdots,\mathbf{o}_{n})$ (17f)
$\mathbf{o}_{concat\_linear}=W\cdot\mathbf{o}_{concat}+b$ (17g)
Figure 3: Various data combinations are performed using element-wise operations (additions $+$ and multiplications $\odot$) over two given outputs. The outputs always have the same data dimensions as the inputs.
Figure 3 presents data combinations over two given function outputs $\mathbf{o_{1}}$ and $\mathbf{o_{2}}$ using element-wise operations. The results are 5 combined outputs: sum, sum + prod, prod, quad, and cubic. Output combinations can utilize higher-degree functions, but these may significantly increase computational complexity, especially in matrix multiplication. Additionally, using more functions results in a larger number of outputs, which can further complicate data combination calculations. To manage this complexity, we prefer to restrict output combinations to quadratic functions involving up to two outputs. For instance, to combine DoG and B-Splines at the output, we can use the following quadratic function formula:
@@
$=\cdots+\cdots+\cdots+f_{DoG}(\mathbf{x})\odot f_{DoG}(\mathbf{x})+f_{BS}(\mathbf{x})\odot f_{BS}(\mathbf{x})$